OpenStudy (anonymous):

Help please! :) use the dot product to find the angle (in degrees) between v=(-6,13) and the vector (1,0).

OpenStudy (psymon):

So our formula for this is: \[\cos \theta = \frac{ u*v }{ ||u||*||v|| }\] Of course I mean the numerator to be a dot, but yeah. So now we just need to use the dot product on your two vectors. Do you know how to do that?

OpenStudy (anonymous):

not really :/

OpenStudy (psymon):

Well, you have (-6i, 13j) and (1i, 0j). You need to multiply the i terms and multiply the j terms, then combine them together. Your answer will be a number, so we drop the i and j, I just have them there so you know which terms to multiply.

OpenStudy (anonymous):

-6=i and 0=j so do I put these numbers in the formula?

OpenStudy (psymon):

Well, you would then add them, which just gives you -6. So -6 goes in the numerator and we now have: \[\cos \theta = \frac{ -6 }{ ||u||*||v|| }\] Now we just need to calculate the magnitudes of u and v. Do you know how to find the magnitude?

OpenStudy (anonymous):

it sounds really familiar but I don't know how to find it

OpenStudy (psymon):

Magnitude is the vector terminology for the length of the vector really. So a vector usually is (ai + bj) where i acts like x-coordinates and j acts like y-coordinates. So just for example, let's graph your first vector.

OpenStudy (psymon):

 Would you know how to find the length of that?

OpenStudy (anonymous):

no lol

OpenStudy (psymon):

Distance formula/pythagorean theorem ^_^